Please give the answers as short as possible, except the last one(a calculation), thanks

What is the H0 null hypothesis in standard deviation analysis?
What does the OC (Operation Characteristics) curve show?
Based on 4-element samples determine the value of the ratio: standard deviation of the observations / standard deviation of the sample averages.
In a two-factor ANOVA test, one factor has a = 3 and the other has b = 4, and each "possible cell" has r = 5 outputs. What is the total degree of freedom (n-1)?
Based on our measurements, the settings of a production line changes without doubt. Regarding their magnitude what is the relation between the actual and critical values of the Abbe test?
What does the AQL abbreviation mean?
Describe which test is suitable for the analysis of variance! Define the concepts involved! How do we determine the parameters of a statistical model?
What is the theoretical distribution of quality control methods used to model the occurrence of debris in the sample?

Using the least-squares method a trendline was fitted to the data below. What is the exact value of the determination coefficient?

x_i=1,2,3     y_i= 7, 5, 3
Answer #1

# in terms of standard deviation null hypothesis will be as follows

H0: 1=2

# oc curve displays the probability of acceptance versus percentage of defective items

# ratio = s/(s/√n)

= √n

# df=(a-2)×(b-1) = 2×3= 6

#

#The decision on rejecting the hypothesis on equality of the means is taken if the value of the statistic SA is less than
the critical value .

#acceptable quality limit (AQL)

# t test

# homogenous distribution

please like ?